# Text mining arXiv: a look through

### 03 Jan 2024 v1

@@ 2 Data description
  • trading and market microstruture (q-fin.TR) studies market microstructure, liquidity, exchange and auction design, automated trading, agent-based modeling and market-making.
These categories are assigned by the authors when they submit their papers. Even if it is possible to select multiple couples of domain-category belonging to more than one domain, we select as reference category only the first category within the quantitative finance domain. Figure 1 shows the numbers of papers on quantitative finace submitted to arXiv between 1997 and 2022. The increase in the last three years is mainly due to the q-fin.EC category.
@@ 3 Text preprocessing
This section describes the text processing steps. The preprocessing phase is performed with nltk:(1) we split the text in tokens; (2) we extract the numbers representing years in the text;333 We assume these numbers have 4 digits. We do not explore this data in the empirical analysis. (3) we identify all strings containing alphabet letters, and we refer to them as words even in the case they do not belong to the English vocabulary; this step allows also to remove some symbols which are not recognized as letters in text analysis. Then, (4) we remove all stopwords and all words with lenght less than 3 characters; we also check whether there are words with more than 25 characters (quite uncommon in English); (5) we conduct a lemmatization by means of a part-of-speech tagger considering nouns, verbs, adjectives and adverbs; (6) we check if the paper is written in English by means of the langdetect package and we discard all non English papers. Both the extraction phase and the preliminary text analysis is parallelized by means of the multiprocessing package. We refer to the output of this first preprocessing phase as lemmatized data.
@@
After a first preprocessing phase, we conduct an $n$-gram analysis by considering the Phrases model of the gensim package. We ignore all words and bigrams with total collected count over the entire corpus lower than 250 and set the score threshold equal to 10. We find the bigrams and then, to find trigrams and fourgrams, we apply again the same model to the trasformed corpus including bigrams. This approach allows us to have a better ex-post understanding of the corpus which is full of $n$-grams (e.g. Monte Carlo simulation, Eisenberg and Noe, or bank balance sheet). The wordclouds of the main bigrams and trigrams (fourgrams) are depicted in Figure 4.
It should be noted that some topic modeling algorithms analized in Section 4 do not need text preprocessing. In those cases the input is just a single list containing the whole paper text. While we refer to this latter input as raw data, we define the output of the preprocessing as cleaned data.
@@ 4.1 Algorithms performance on full texts
In theory both BERTopic and Top2Vec should use raw data since these algorithms rely on an embedding approach, and keeping the original structure of the text is of paramount importance (see Egger and Yu (2022)). However, raw data extracted from quantitative finance papers have a considerable amounts of formulas, symbols and numbers that may affect the algorithms performance. For this reason, we consider both raw and cleaned data. Additionally, for these state-of-the-art algorithms the number of extracted topics tends be large, but the algorithms offer the possibility to reduce the number of topics and of outliers, that can be larger than expected. The parameters of a BERTopic model have to be carefully chosen to avoid memory issues. Alternativelly, it is possible to perform an online topic modeling, that is the model is trained incrementally from a mini-batch of instances. This result in a less resource demanding approach in terms of memory and CPU usage, but it generates less rich and less comprehensive outputs and for these reasons we do not consider this incremental approach here.
@@ 4.2 Empirical study
As shown in Section 4.1, the best performing model is Doc2Vec with K-means clustering applied on cleaned data. This model is considered to have a better understanding of the topics discussed in the quantitative finance papers analized in this work. To obtain the desired number of topics we perform again a K-means clustering analysis. To extract the most representative documents, we retrieve the centroid vectors of each cluster. These centroids represent the average position of all document embeddings assigned to a particular cluster. For each cluster centroid, we find the nearest neighbors among the original Doc2Vec embeddings. These nearest neighbors are the documents that are closest to the centroid in the embedding space and can be considered as the main documents of that cluster. Thus we select for each cluster the 20 most representative documents and we find a label for the topic on the basis of the documents titles. Note that the underlying meanings of the topics are subject to human interpretation. However, also this phase is automated by asking the topic to ChatGPT (GPT-3.5) after having provided the list of 20 titles (see Ebinezer (2023)).
@@
It is interesting to observe that topics related to decentralized finance and blockchain technology (2) and stock price prediction with deep learning and news sentiment analysis (20) show a remarkable increase. This is also true for health, policy, and social impact studies, represented by the topic 16, and diverse perspectives in education, innovation, and economic development (0). Both topics are oriented towards economics. These topics trends depend also by the introduction in 2014 of the q-fin.EC category within the arXiv quantitative finance papers. Classical quantitative finance subjects like portfolio optimization techniques and strategies (6), stochastic volatility modeling and option pricing (7), game theory and strategic decision-making (19) high-order numerical methods for option pricing in finance (23) as well as, new themes appeared in the literature in the last years, like deep reinforcement learning in stock trading and portfolio management (4) and environmental and economic impacts of mobility technologies (25) have attracted the interest of researchers in the analized period. The representativeness of topic 28 is limited, mainly because the number of papers in this cluster is low, and one could merge it with another cluster (i.e. 8).
@@ 5 Extracting authors and journals
In this section we extract the author surnames by means of the spacy package. More in details, starting from the raw data, we perform a named-entities recognition. Since this approach extract both first names and surnames, we remove all first names by checking if these names are included in a list of about 67,000 first names. It should be noted that the number of occurence of the surname of the author in a paper strongly depends on the citation style. Surnames are always reported in the references, but they do not necessary appear in the main text of a paper. Additionally, even if the author-date style is widely used (i.e. the citation in the text consists of the authors name and year of publication), the surname of the first author appears more frequently (e.g. Bianchi is more probable than Tassinari, even if Bianchi and Tassinari are coauthors of the same papers, together with other coauthors).
@@
The algorithm is able to find the names and surnames occuring in the text. These are included in the PERSON entity types. The first 100 authors by number of occurences in the corpus are selected. In order to have additional information on these authors, we obtain topics, number of citations, h-index and i10-index from Google Scholar (see Table 3). It should be noted that not all authors are registered in Google Scholar, even if they made a significant contribution to the field (e.g. Markowitz) or there are authors with the same surname and belonging to the same research field (see also Figure 3 in Sharma et al. (2023)). This is the case for some researchers we find in our corpus (e.g. Zhou, Bayraktar and Chakrabarti). For these last authors is not simple to find a perfect match in Google Scholar even if their number of occurrences is generally high.444 For the reasons described above we exclude from Table 3 the following researcher: Zhou, Markowitz, Peng, Jacod, Merton, Guo, Lo, Follmer, Yor, Almgren, Embrechts, Bayraktar, Artzner, Weber, Jarrow, Feng, Samuelson, Tang, Chakrabarti, Glasserman, Tsallis, Leung, Sato, Zariphopoulou, Kramkov, Karoui, Cizeau, Cao and Christensen.
@@
The algorithm is also able to find the most cited journals, included in the ORG entity types: Journal of Finance (4490 $*{}^{*}$, Mathematical Finance (3785), Journal of Financial Economics ($*{}^{*}$, Physica A ($*{}^{*}$, Quantitative Finance (3044), Econometrica ($*{}^{*}$, Journal of Econometrics (1971), American Economic Review (1878), Insurance Mathematics and Economics (1667), Review of Financial Studies ($*{}^{*}$, Journal of Banking and Finance ($*{}^{*}$, Physical Review (1538), Journal of Economic Dynamics (1289), Energy (1267), Operations Research (1242), The Quarterly Journal of Economics ($*{}^{*}$, Journal of the American Statistical Association (1204), Management Science (1160), European Journal of Operational Research (1066), Quantum (1043), IEEE Transactions (996), Journal of Political Economy (990), Journal of Economic Theory (977), Energy Economics (946), International Journal of Theoretical and Applied Finance (946), SIAM Journal on Financial Mathematics (888), Science (865), Expert Systems with Applications (845), Applied Mathematical Finance (743), Finance and Stochastics (736), Mathematics of Operations Research (652), PLoS (602), The Annals of Applied Probability (593), Stochastic Processes and their Applications (525), Energy Policy (520), Internationl Journal of Forecasting (520), The European Physical Journal (514), Journal of Empirical Finance (510), Journal of Risk (509). We consider only journals with more than 500 occurrences and we esclude publishing houses. The journals with the asterisk symbol are identified with more than one name. It should be noted that some well-known journals are slightly below 500 occurences. Furthermore, it is worth noting that papers with a strong mathematical focus tend to receive significantly fewer citations compared to papers that lean more towards economics or finance.

### 05 Apr 2024 v2

@@ 2 Data description
  • trading and market microstructure (q-fin.TR) studies market microstructure, liquidity, exchange and auction design, automated trading, agent-based modeling and market-making.
These categories are assigned by the authors when they submit their papers. Even if it is possible to select multiple couples of domain-category belonging to more than one domain, we select as reference category only the first category within the quantitative finance domain. Figure 1 shows the numbers of papers on quantitative finance submitted to arXiv between 1997 and 2022. The increase in the last three years is mainly due to the q-fin.EC category.
@@ 3 Text preprocessing
This section describes the text processing steps. The preprocessing phase is performed with nltk:(1) we split the text in tokens; (2) we extract the numbers representing years in the text;333 We assume these numbers have 4 digits. We do not explore this data in the empirical analysis. (3) we identify all strings containing alphabet letters, and we refer to them as words even in the case they do not belong to the English vocabulary; this step allows also to remove some symbols which are not recognized as letters in text analysis. Then, (4) we remove all stopwords and all words with length less than 3 characters; we also check whether there are words with more than 25 characters (quite uncommon in English); (5) we conduct a lemmatization by means of a part-of-speech tagger considering nouns, verbs, adjectives and adverbs; (6) we check if the paper is written in English by means of the langdetect package and we discard all non English papers. Both the extraction phase and the preliminary text analysis is parallelized by means of the multiprocessing package. We refer to the output of this first preprocessing phase as lemmatized data.
@@
After a first preprocessing phase, we conduct an $n$-gram analysis by considering the Phrases model of the gensim package. We ignore all words and bigrams with total collected count over the entire corpus lower than 250 and set the score threshold equal to 10. We find the bigrams and then, to find trigrams and fourgrams, we apply again the same model to the transformed corpus including bigrams. This approach allows us to have a better ex-post understanding of the corpus which is full of $n$-grams (e.g. Monte Carlo simulation, Eisenberg and Noe, or bank balance sheet). The wordclouds of the main bigrams and trigrams (fourgrams) are depicted in Figure 4.
It should be noted that some topic modeling algorithms analyzed in Section 4 do not need text preprocessing. In those cases the input is just a single list containing the whole paper text. While we refer to this latter input as raw data, we define the output of the preprocessing as cleaned data.
@@ 4.1 Algorithms performance on full texts
In theory both BERTopic and Top2Vec should use raw data since these algorithms rely on an embedding approach, and keeping the original structure of the text is of paramount importance (see Egger and Yu (2022)). However, raw data extracted from quantitative finance papers have a considerable amounts of formulas, symbols and numbers that may affect the algorithms performance. For this reason, we consider both raw and cleaned data. Additionally, for these state-of-the-art algorithms the number of extracted topics tends be large, but the algorithms offer the possibility to reduce the number of topics and of outliers, that can be larger than expected. The parameters of a BERTopic model have to be carefully chosen to avoid memory issues. Alternatively, it is possible to perform an online topic modeling, that is the model is trained incrementally from a mini-batch of instances. This result in a less resource demanding approach in terms of memory and CPU usage, but it generates less rich and less comprehensive outputs and for these reasons we do not consider this incremental approach here.
@@ 4.2 Empirical study
As shown in Section 4.1, the best performing model is Doc2Vec with K-means clustering applied on cleaned data. This model is considered to have a better understanding of the topics discussed in the quantitative finance papers analyzed in this work. To obtain the desired number of topics we perform again a K-means clustering analysis. To extract the most representative documents, we retrieve the centroid vectors of each cluster. These centroids represent the average position of all document embeddings assigned to a particular cluster. For each cluster centroid, we find the nearest neighbors among the original Doc2Vec embeddings. These nearest neighbors are the documents that are closest to the centroid in the embedding space and can be considered as the main documents of that cluster. Thus we select for each cluster the 20 most representative documents and we find a label for the topic on the basis of the documents titles. Note that the underlying meanings of the topics are subject to human interpretation. However, also this phase is automated by asking the topic to ChatGPT (GPT-3.5) after having provided the list of 20 titles (see Ebinezer (2023)).
@@
It is interesting to observe that topics related to decentralized finance and blockchain technology (2) and stock price prediction with deep learning and news sentiment analysis (20) show a remarkable increase. This is also true for health, policy, and social impact studies, represented by the topic 16, and diverse perspectives in education, innovation, and economic development (0). Both topics are oriented towards economics. These topics trends depend also by the introduction in 2014 of the q-fin.EC category within the arXiv quantitative finance papers. Classical quantitative finance subjects like portfolio optimization techniques and strategies (6), stochastic volatility modeling and option pricing (7), game theory and strategic decision-making (19) high-order numerical methods for option pricing in finance (23) as well as, new themes appeared in the literature in the last years, like deep reinforcement learning in stock trading and portfolio management (4) and environmental and economic impacts of mobility technologies (25) have attracted the interest of researchers in the analyzed period. The representativeness of topic 28 is limited, mainly because the number of papers in this cluster is low, and one could merge it with another cluster (i.e. 8).
@@ 5 Extracting authors and journals
In this section we extract the author surnames by means of the spacy package. More in details, starting from the raw data, we perform a named-entities recognition. Since this approach extract both first names and surnames, we remove all first names by checking if these names are included in a list of about 67,000 first names. It should be noted that the number of occurrence of the surname of the author in a paper strongly depends on the citation style. Surnames are always reported in the references, but they do not necessary appear in the main text of a paper. Additionally, even if the author-date style is widely used (i.e. the citation in the text consists of the authors name and year of publication), the surname of the first author appears more frequently (e.g. Bianchi is more probable than Tassinari, even if Bianchi and Tassinari are coauthors of the same papers, together with other coauthors).
@@
The algorithm is able to find the names and surnames occurring in the text. These are included in the PERSON entity types. The first 100 authors by number of occurrences in the corpus are selected. In order to have additional information on these authors, we obtain topics, number of citations, h-index and i10-index from Google Scholar (see Table 3). It should be noted that not all authors are registered in Google Scholar, even if they made a significant contribution to the field (e.g. Markowitz) or there are authors with the same surname and belonging to the same research field (see also Figure 3 in Sharma et al. (2024)). This is the case for some researchers we find in our corpus (e.g. Zhou, Bayraktar and Chakrabarti). For these last authors is not simple to find a perfect match in Google Scholar even if their number of occurrences is generally high.444 For the reasons described above we exclude from Table 3 the following researcher: Zhou, Markowitz, Peng, Jacod, Merton, Guo, Lo, Follmer, Yor, Almgren, Embrechts, Bayraktar, Artzner, Weber, Jarrow, Feng, Samuelson, Tang, Chakrabarti, Glasserman, Tsallis, Leung, Sato, Zariphopoulou, Kramkov, Karoui, Cizeau, Cao and Christensen.
@@
The algorithm is also able to find the most cited journals, included in the ORG entity types: Journal of Finance (4490 $*{}^{*}$, Mathematical Finance (3785), Journal of Financial Economics ($*{}^{*}$, Physica A ($*{}^{*}$, Quantitative Finance (3044), Econometrica ($*{}^{*}$, Journal of Econometrics (1971), American Economic Review (1878), Insurance Mathematics and Economics (1667), Review of Financial Studies ($*{}^{*}$, Journal of Banking and Finance ($*{}^{*}$, Physical Review (1538), Journal of Economic Dynamics (1289), Energy (1267), Operations Research (1242), The Quarterly Journal of Economics ($*{}^{*}$, Journal of the American Statistical Association (1204), Management Science (1160), European Journal of Operational Research (1066), Quantum (1043), IEEE Transactions (996), Journal of Political Economy (990), Journal of Economic Theory (977), Energy Economics (946), International Journal of Theoretical and Applied Finance (946), SIAM Journal on Financial Mathematics (888), Science (865), Expert Systems with Applications (845), Applied Mathematical Finance (743), Finance and Stochastics (736), Mathematics of Operations Research (652), PLoS (602), The Annals of Applied Probability (593), Stochastic Processes and their Applications (525), Energy Policy (520), International Journal of Forecasting (520), The European Physical Journal (514), Journal of Empirical Finance (510), Journal of Risk (509). We consider only journals with more than 500 occurrences and we exclude publishing houses. The journals with the asterisk symbol are identified with more than one name. It should be noted that some well-known journals are slightly below 500 occurrences. Furthermore, it is worth noting that papers with a strong mathematical focus tend to receive significantly fewer citations compared to papers that lean more towards economics or finance.